OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 A. 0 B. 4 C. x can be any number as there are infinitely many solutions to the system. D. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

D [THERE IS NO X(MEANING THE SYSTEM IS INCONSISTENCE)]

OpenStudy (anonymous):

if u double the first equation u can see it is the same as the second. So there are said to be infinitely many solutions in this case. Then the answer is C

OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations auditioned today?

OpenStudy (anonymous):

x+y=1150 3x-2y = 1300 U should be able to do it by now, double first one and add...

OpenStudy (anonymous):

-y = -2x+6 y = 2x-6 4x-2(2x-6)=12 4x-4x+12=12 0=0 D

OpenStudy (anonymous):

Another one ..... If the system reduces to 0=0 (true) then the system is said to have an infinite number of solutions.

OpenStudy (anonymous):

He's right.